three equal weight A,B,C of mass 2 kg each are hang8ng on a string passing over a fixed frictionless pulley the tension in the string connecting weightsB and C

Grade:11

1 Answers

Aman gohel
6 years ago
Given,A=B=C=4KgThe total weight on string=B+C=2Kg+2kg=4Kg=>Therefore tension on string=m×g=4×10=40N(Ans)